Suppose that every year the city of New York auctions off a limited number of medallions that give cab drivers an exclusive right to taxi in the city for one year. After receiving the medallion the cab driver is essentially a monopolist facing an inverse demand function p = 20 - 1.8Q. The total cost function for each cab driver can be summarized as C(Q) = 3Q + F. Where F represents the cost of the medallion or annual fixed cost. What would each taxi medallion have to cost in the long run for this to be a case of monopolistic competition.
Answer #1

Marginal cost (MC) = dC(Q)/dQ = 3

Average cost (ATC) = C(Q)/Q = 3 + (F/Q)

In the long run equilibrium of monopolistic competition, MR = MC and P = ATC.

Total revenue (TR) = P x Q = 20Q - 1.8Q2

MR = dTR/dQ = 20 - 3.6Q

Equating with MC,

20 - 3.6Q = 3

3.6Q = 17

Q = 4.72

P = 20 - (1.8 x 4.72) = 20 - 8.5 = 11.5

ATC = 3 + (F/4.72)

Equating P and ATC,

11.5 = 3 + (F/4.72)

F/4.72 = 8.5

F = 40.12